### v1

# Acceleration of the CASINO quantum Monte Carlo software using graphics processing units and OpenACC

Abstract
We describe how quantum Monte Carlo calculations using the CASINO software can be accelerated using graphics processing units (GPUs) and OpenACC. In particular we consider offloading Ewald summation, the evaluation of long-range two-body terms in the Jastrow correlation factor, and the evaluation of orbitals in a blip basis set. We present results for three- and two-dimensional homogeneous electron gases and ab initio simulations of bulk materials, showing that significant speedups of up to a factor of 2.5 can be achieved by the use of GPUs when several hundred particles are included in the simulations. The use of single-precision arithmetic can improve the speedup further without significant detriment to the accuracy of the calculations.
keywords:
Quantum Monte Carlo , Graphics processing units , OpenACC , Fortran
journal: Computer Physics Communications
1 Introduction
Continuum quantum Monte Carlo (QMC) methods are powerful techniques for finding approximate solutions to the many-body Schrödinger equation. In these methods the quantum many-body problem is tackled directly in the continuous, $3N$-dimensional position basis, with Monte Carlo methods being used to explore and integrate over the high-dimensional configuration space. In this article we focus on the variational and diffusion quantum Monte Carlo (VMC and DMC) methods, which allow one to find the ground state and selected excited states of fermionic and bosonic systems, as implemented in the CASINO software [1]. A major challenge for these methods is the computational expense, which not only constrains the range of systems that can be studied, but also limits the accuracy that can be achieved, as users are forced to use smaller simulation supercells, larger time steps, etc. The high cost of QMC methods is partially offset by the fact that QMC algorithms are inherently parallelizable: one can reduce error bars on estimated quantities by averaging over independent random walks. Hence, provided equilibration time is a negligible fraction of total run time, QMC methods are almost perfect for distributed memory message-passing interface (MPI) parallelization, and CASINO has been shown to scale to tens of thousands of MPI processes [2].
In addition to distributed memory parallelism, CASINO allows OpenMP shared memory parallelism in loops over particles in calculations of the potential energy and calculations of various parts of the many-body wave function. This allows one to reduce the walltime for the equilibration stage of a QMC calculation, and is generally highly efficient when there are hundreds of particles. However, efficient parallelism on CPUs is not the only consideration when examining whether software is suitable for use on modern high-performance computers; it is also important to assess whether software can use accelerators such as graphics processing units (GPUs). This is the issue that we address in this article.
Since around 2010 the Top 500 list [3] has been increasingly dominated by supercomputers featuring GPU accelerators. At the time of writing the most powerful computer in the world is Frontier at Oak Ridge National Laboratory, which was the first exascale computer. Each compute node of Frontier has one 64-core AMD Epyc 7713 “Trento” CPU and four 220-core AMD Instinct MI250X GPUs; hence Frontier is heavily reliant on GPUs to achieve its peak performance. The growth of machine learning has both driven and been enabled by the availability of GPU accelerators for training neural networks. The need to reduce the energy consumption per floating-point operation has provided a further stimulus to the rise of GPUs in high-performance computing.
One of the major players in the supercomputing GPU market, the NVIDIA Corporation, has developed the proprietary compute unified device architecture (CUDA) platform and application programming interface (API) [4]. CUDA allows programmers to have full control over the offloading of work to NVIDIA GPUs, but is not compatible with the hardware or compilers of other vendors. In practice, were a Fortran program such as CASINO to adopt CUDA, it would be necessary to fork the software into CUDA and non-CUDA versions, which would pose problems of maintainability and portability. Despite the fact that our work focuses on NVIDIA hardware, we are keen to avoid these challenges associated with the use of CUDA, and we have adopted solutions that allow us to retain a single code base as well as the support of a plurality of compiler and hardware vendors.
A directives-based approach provides a viable technique for allowing control of offloading from within a single Fortran code base. The OpenACC programming standard and API [5] is supported by Fortran compilers including NVIDIA’s nvfortran, the Cray/HPE Fortran compiler, and GNU’s gfortran. OpenACC is developed and overseen by the OpenACC Organization, a non-profit organization whose members include the NVIDIA Corporation, AMD, HPE, ARM, and Oak Ridge National Laboratory (home to the Frontier supercomputer), as well as many HPC centers and academic institutions. OpenACC compiler directives are used to identify parts of a program that should be compiled for execution on an accelerator device; furthermore these directives are also used to provide much of the functionality specific to programming for GPUs.
These directives include:
  • DATA directives, to control the transfer of data between the main memory and the device memory.
  • ASYNC and WAIT directives, to allow asynchronous kernel operations for overlapping computation with data movement.
  • GANG, WORKER, and VECTOR directives, to enable developers to map multidimensional arrays and nested loops more directly to the available hardware without relying on the compiler.
A second directives-based approach is the extension of the OpenMP API [6] to support offloading to generic accelerator devices, including GPUs. OpenMP is developed by the OpenMP Architecture Review Board, which is supported by a broad range of vendors in the high-performance computing sector. OpenMP provides a well-established and widely used standard for handling multiple CPU threads in numerical software. As with OpenACC, offloading to accelerators is controlled by means of compiler directives, to compile specific sections of the program for execution on an accelerator. These also allow the developer to control data movement between the device and the host CPU, as well as performing asynchronous operations and a limited amount of nested/multilevel parallelism through the use of single-instruction, multiple-data (SIMD) teams.
From a developer’s point of view, the existence of similar, but incompatible, competing standards, each with varying levels of support from the different vendors, is not ideal. In this work we have chosen to use OpenACC, primarily because the NVIDIA developer ecosystem is the most mature, and currently has better support for the OpenACC API than OpenMP. Furthermore, OpenACC is specifically designed with accelerators in mind and allows the programmer and the compiler more scope to optimize the directives to make best use of the underlying hardware. If in the future OpenMP should emerge as the dominant standard, it will be relatively easy for us to adapt our OpenACC implementation to OpenMP, as OpenMP is already in use within CASINO for CPU parallelism.
Other QMC codes have each taken a different approach for supporting GPU computing. In the early 2010s, the QMCPACK code [7] forked into a limited-feature CUDA version and a full-feature non-CUDA version, but QMCPACK has more recently adopted OpenMP for offloading from a single code base. The PyQMC module [8] in PySCF [9] makes use of the CuPy library [10], which in turn uses CUDA and AMD’s ROCm software stack for GPU computing [11], to allow the use of GPUs to complete various tasks. The European version of the CHAMP code [12] can make use of NVIDIA’s cuBLAS [13] library for linear algebra, while the North American version of CHAMP [14] supports GPUs via the einspline library for spline representation of parts of the wave function [15]. The TurboRVB QMC code uses OpenMP for offloading [16]. To our knowledge, apart from an early experiment [17] with the QMcBeaver code [18], OpenACC has not been used in a major QMC code.
The rest of this paper is structured as follows. In Section 2 we describe the aspects of a CASINO calculation that are particularly promising for offloading to GPUs. The tools used for profiling CASINO on CPUs and GPUs are discussed in Section 3. In Section 4 we take a small but useful diversion into how exactly one can quantify GPU performance. In Section 5 we describe our modifications to CASINO to allow the use of GPUs. The use of single precision (SP) arithmetic rather than double precision (DP) arithmetic is discussed in Section 6. We present performance results in Section 8. Finally, we draw our conclusions in Section 9. Throughout we use Hartree atomic units ($\hbar=m_{\rm e}=|e|=4\pi\epsilon_{0}=1$ a.u.), except where otherwise stated.
2 Targets in CASINO for offloading to GPUs
2.1 VMC and DMC methods
In the VMC method we use the Metropolis algorithm to sample particle configurations distributed as the square modulus of a trial spatial wave function $\Psi$, and we average $(\hat{A}\Psi)/\Psi$ over these configurations to obtain an unbiased estimate of $\langle\Psi|\hat{A}|\Psi\rangle/\langle\Psi|\Psi\rangle$, where $\hat{A}$ is a spin-independent quantum operator [19]. Free parameters in the trial wave function are optimized by minimizing either the energy expectation value or the variance of the energy.
In the DMC method [19, 20] the ground-state component of the trial wave function is projected out by simulating drift, diffusion, and branching/dying processes governed by the time-dependent Schrödinger equation in imaginary time. For the ground state of a bosonic system there are no uncontrolled approximations; for fermionic systems or for excited states we solve the imaginary-time-dependent Schrödinger subject to the boundary condition that the wave function remains fixed at zero at the nodes of the trial wave function $\Psi$. Expectation values are computed in a similar manner to VMC.
In practice the VMC and DMC methods both involve repeatedly proposing single-particle moves, calculating the value and first and second derivatives of the trial wave function $\Psi$, and calculating (changes in) the potential energy.
The trial wave functions used in QMC calculations are typically of Slater-Jastrow-backflow form
$\Psi({\bf R})=e^{J({\bf R})}S({\bf X}({\bf R})),$ (1)
where $S({\bf R})$ is an expansion of one or more Slater determinants of single-particle orbitals, $e^{J({\bf R})}$ is a Jastrow correlation factor and ${\bf X}({\bf R})$ is a backflow transformation of the $3N$-dimensional vector of particle coordinates ${\bf R}$. The orbitals in the Slater determinant are usually taken from either density functional theory (DFT) calculations or wave-function-based quantum chemistry calculations, e.g., Hartree-Fock theory. In this work we will focus on single-determinant Slater-Jastrow wave functions without a backflow transformation [i.e., we will assume that ${\bf X}({\bf R})={\bf R}$ and that $S$ is the product of a Slater determinant of occupied orbitals for spin-up electrons and a Slater determinant of occupied orbitals for spin-down electrons]. CASINO’s standard Jastrow factor [21] is described in Section 2.3.
Each time step of a Slater-Jastrow QMC simulation of $N$ particles requires the evaluation of $O(N^{2})$ long-range pairwise Jastrow terms, each of $O(N)$ orbitals expanded in a localized basis set at each of the $N$ particle positions, and the sum of the $O(N^{2})$ pairwise Coulomb interactions; the cost hence scales as $O(N^{2})$. The cost of updating the Slater determinants is $O(N^{3})$ but with a sufficiently small prefactor that this is only relevant at the very largest system sizes that have been studied using QMC (1000–2000 particles). As an aside, we note that the number of time steps required to achieve a given statistical error bar must grow as the system size to counteract the growth in the variance, and hence the scaling of standard Slater-Jastrow QMC is $O(N^{3}+\epsilon N^{4})$ [19].
2.2 Ewald summation
In a molecular system, the electron-electron Coulomb potential energy is simply a sum of pairwise $1/r$ interactions. Even with thousands of electrons, this is cheap to evaluate. On the other hand, to model bulk condensed matter, we study simulation supercells subject to periodic boundary conditions and hence we need to use periodic Coulomb potentials such that the electrostatic energy tends to a well-defined and appropriate value in the limit of infinite simulation-cell size.
The most well-known and widely used method for evaluating Coulomb interactions in periodic (crystalline) systems is Ewald summation [22]. The Coulomb potential is taken to be the periodic solution to Poisson’s equation, which is evaluated by means of rapidly convergent series in real space and reciprocal space. Physically, the use of the periodic solution of Poisson’s equation corresponds to the absence of a macroscopic polarization field throughout the crystal, as would be the case if the macroscopic crystal were wrapped in a perfect conductor so that surface polarization charges were cancelled by image charges in the conductor. The electrostatic energy of a set of charges $\{q_{i}\}$ with positions $\{{\bf r}_{i}\}$ in a periodic simulation cell is
$\hat{V}_{\rm Ew}=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}q_{i}q_{j}v_{\rm E}({\bf r}_{i}-{\bf r}_{j})+\frac{1}{2}\sum_{j}q_{j}^{2}v_{\rm M},$ (2)
where
$v_{\rm E}({\bf r})$ $=$ $\frac{4\pi}{\Omega_{\rm s}}\sum_{{\bf G}_{\rm s}\neq{\bf 0}}\frac{e^{-G_{\rm s}^{2}/(4\gamma)+i{\bf G}_{\rm s}\cdot{\bf r}}}{G_{\rm s}^{2}}+\sum_{{\bf R}_{\rm s}}\frac{\operatorname{erfc}\left(\sqrt{\gamma}|{\bf r}-{\bf R}_{\rm s}|\right)}{|{\bf r}-{\bf R}_{\rm s}|}-\frac{\pi}{\Omega_{\rm s}\gamma}$ (3)
$v_{\rm M}$ $=$ $\frac{4\pi}{\Omega_{\rm s}}\sum_{{\bf G}_{\rm s}\neq{\bf 0}}\frac{e^{-G_{\rm s}^{2}/(4\gamma)}}{G_{\rm s}^{2}}+\sum_{{\bf R}_{\rm s}\neq{\bf 0}}\frac{\operatorname{erfc}\left(\sqrt{\gamma}R_{\rm s}\right)}{R_{\rm s}}-2\sqrt{\frac{\gamma}{\pi}}-\frac{\pi}{\Omega_{\rm s}\gamma},$ (4)
and $\Omega_{\rm s}$ is the volume of the simulation cell, $\{{\bf R}_{\rm s}\}$ are the simulation-cell lattice points, $\{{\bf G}_{\rm s}\}$ are the simulation-cell reciprocal lattice points, and $\gamma$ is a constant that can be chosen to minimize the cost of evaluating Eq. (3) [22]. Similar expressions can be derived for Ewald interactions in two dimensions [23, 24, 25].
It is clear that the cost of evaluating the Ewald interaction in Eq. (3) is vastly greater than the cost of evaluating $1/r$. Equation (3) is suitable for offloading to GPUs because the evaluation of the series is dominated by floating point operations and both the real- and reciprocal-space series are inherently parallelizable; moreover, much of the data required can be moved to the GPU once, at the outset of the calculation. Offloading of Eq. (3) is investigated in Section 5.1.
2.3 Long-range Jastrow terms
CASINO’s standard Jastrow exponent is of the form
$J$ $=$ $\sum_{i=1}^{N}\left[\sum_{I=1}^{N_{\rm n}}\chi(r_{iI})+q({\bf r}_{i})\right]+\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\left[u(r_{ij})+p({\bf r}_{ij})\right]$ (5)
${}+\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\sum_{I=1}^{N_{\rm n}}f(r_{iI},r_{jI},r_{ij})+\sum_{i=1}^{N-2}\sum_{j=i+1}^{N-1}\sum_{k=j+1}^{N}H(r_{jk},r_{ki},r_{ij}),$
where $\chi(r)$, $u(r)$, $f(r,r^{\prime},r^{\prime\prime})$, and $H(r,r^{\prime},r^{\prime\prime})$ are polynomial functions of interparticle distances, and $p({\bf r})$ and $q({\bf r})$ are plane-wave expansions in the reciprocal lattice points of the simulation cell and the primitive cell, respectively [21, 26]. $N_{\rm n}$ is the number of nuclei, ${\bf r}_{i}$ is the position of particle $i$, ${\bf r}_{ij}$ is the nearest-image displacement between particles $i$ and $j$, and $r_{iI}$ is the nearest-image distance between particle $i$ and nucleus $I$. The functions $u$, $\chi$, etc., may depend on the species of particle involved (e.g., $u$ is nearly always different for parallel- and antiparallel-spin electrons due to differences in the Kato cusp conditions) and, where relevant, the functions depend on the ions involved.
The four terms in Eq. (5) are, respectively, an inhomogeneous one-body term, a homogeneous two-body term, an inhomogeneous two-body term, and a homogeneous three-body term. The Jastrow factor does not affect the wave-function (anti)symmetry described by the Slater part of the wave function. The Jastrow factor is nodeless and hence allows a description of correlation without altering the excitation level described by the Slater wave function. The polynomial two-body term $u(r)$ is used to enforce the electron-electron Kato cusp conditions [27], and the polynomial one-body term $\chi(r)$ can be used to enforce the electron-nucleus cusp conditions. The other terms are required to be at least twice differentiable everywhere.
The homogeneous two-body correlations $u$ and $p$ describe both short-range dynamical correlations (the Kato cusp behavior) and long-range behavior that is well-described by the random phase approximation [19]. The inhomogeneous one-body terms $\chi$ and $q$ are primarily required to undo the effects of two-body correlations on the charge density around atoms described by the Slater wave function; hence they are mostly short-ranged. The inhomogeneous two-body correlations $f$ are intended to describe the density dependence of two-body correlations inside atoms, and hence are primarily short-ranged. The homogeneous three-body term $H$ is rarely used except in model systems such as electron-hole complexes [28] or cold atomic gases [29]. Hence the cost of evaluating the Jastrow factor is usually dominated by the evaluation of the homogeneous two-body terms. We therefore focus on offloading the evaluation of the values and derivatives of the two-body terms $u(r)$ and $p({\bf r})$.
The forms of $u$ and $p$ are as follows [21]:
$u(r)$ $=$ $\sum_{l=0}^{N_{u}}\alpha_{l}r^{l}(r-L_{u})^{C}\Theta(L_{u}-r)$ (6)
$p({\bf r})$ $=$ $\sum_{m=1}^{N_{p}}a_{m}\sum_{{\bf G}_{\rm s}\in\bigstar_{m}^{+}}\cos\left({\bf G}_{\rm s}\cdot{\bf r}\right),$ (7)
[…]
For ${\bf G}_{\rm s}=n_{1}{\bf b}_{1}+n_{2}{\bf b}_{2}+n_{3}{\bf b}_{3}$, where $\{{\bf b}_{i}\}$ are supercell reciprocal lattice vectors, we may write
$\cos({\bf G}_{\rm s}\cdot{\bf r}_{ij})=\operatorname{Re}\left[{(e^{i{\bf b}_{1}\cdot{\bf r}_{ij}})}^{n_{1}}\,{(e^{i{\bf b}_{2}\cdot{\bf r}_{ij}})}^{n_{2}}\,{(e^{i{\bf b}_{3}\cdot{\bf r}_{ij}})}^{n_{3}}\right],$ (8)
so only three complex exponentials are required to compute all the required cosines for Eq. (7). The powers of these exponentials are computed and buffered. The required number of reciprocal lattice points is typically $\sim$10$\sim$100. We note that for large plane-wave expansions there is scope to reduce the computational cost of the $p$ term by means of a blip (B-spline) re-representation.
The modifications to CASINO to enable offloading of the long-range two body Jastrow terms are described in Section 5.2.
2.4 Evaluation of orbitals in a blip basis set
Especially in studies of condensed matter, it is common practice to use a plane-wave basis set to represent single-particle orbitals in DFT calculations. However, QMC simulations are performed in real space, and hence a plane-wave basis set is a particularly expensive choice: after each single-particle move, all $O(N)$ plane-wave basis functions must be evaluated at the particle’s new position; evaluation of the $O(N)$ occupied orbitals at the new position requires $O(N^{2})$ multiplications of basis functions by orbital coefficients. To reduce the cost we may re-represent the orbitals in a localized blip (B-spline) basis set before the start of a QMC calculation [30, 31]. The cost of evaluating the orbitals after each particle move is then reduced to the cost of looking up the particle’s position in the spline grid followed by the evaluation of a small number of cubic polynomials defining the blip basis functions, followed by multiplication of the blip functions by coefficients for each orbital [$O(N)$ multiplications].
We investigate offloading the multiplication of the blip basis functions by the blip coefficients for each orbital in Section 5.3. A similar approach could be taken for evaluating orbitals in other basis sets, e.g., Gaussian basis sets.
3 Profile-driven GPU porting of CASINO
3.1 Development cycle
The development cycle for porting CASINO to GPUs using OpenACC is outlined as follows.
  1. (1)
    Profile the application to identify performance bottlenecks.
  2. (2)
    Assess which of the bottlenecks identified in (1) are amenable to acceleration by GPUs; refactor CPU code as necessary to expose parallelism maximally.
  3. (3)
    Use OpenACC to parallelize (or to offload to an appropriate library) a bottleneck prioritized from those in (2).
  4. (4)
    Minimize host-device data migrations resulting from (3).
  5. (5)
    Optimize the directives introduced in (3) and (4) for given architectures and/or inputs.
The cycle is executed incrementally to ensure correctness, and iteratively to achieve the maximum performance improvement.
3.2 Subroutine timers
Profiling CASINO is made straightforward by an internal timer system, which allows the walltime, CPU time, and number of invocations for different activities to be recorded and reported at the end of a calculation. Extensive profiling of CASINO as part of the European Union’s Performance Optimization and Productivity (POP) program [32] yielded results in good agreement with the results of CASINO’s internal timers.
CASINO supports a very broad range of use cases, from studies of molecules with compact Gaussian or Slater-type-orbital basis sets and multideterminant Slater-Jastrow(-backflow) wave functions, to studies of nanomaterials and condensed matter using blip basis sets and Slater-Jastrow(-backflow) wave functions, to studies of plasmas of particles with a range of different interactions moving in one, two, or three dimensions with Slater-Jastrow(-backflow) or pairing wave functions. As a representative example of the breakdown of timings for a large ab initio condensed matter system, we consider a DMC calculation in a $3\times 3\times 3$ supercell of BaTiO3, with a Slater-Jastrow wave function and a blip basis set for the orbitals. The atomic cores are described by Dirac-Fock pseudopotentials [33, 34], such that the supercell contains 648 electrons. A target population of 512 walkers is used, together with a time step of 0.01 Ha-1. In addition to the total energy, the DMC charge density is accumulated. For the baseline CPU timings and profile, CASINO was compiled using the Intel Fortran compiler and run on a 64-core Intel Ice Lake node. CPU timings broken down by activity are shown in Table 1. Activities not accounted for in Table 1 include: the updating and copying of arrays when particle moves are proposed and accepted, and after branching of walkers; evaluating random numbers; checkpointing; etc. Summing energies over MPI processes is expensive because it is a blocking operation, and the populations of walkers assigned to each process fluctuate due to branching.
Table 1: Percentage of CPU time spent on different activities in a 648-electron BaTiO3 DMC calculation with a Slater-Jastrow trial wave function.
Activity Percentage of CPU time
Evaluating Jastrow $u$ terms 1.7%
Evaluating Jastrow $\chi$ terms 0.5%
Evaluating Jastrow $f$ terms 10.4%
Evaluating Jastrow $p$ terms 51.5%
Evaluating e-e and e-n distances 1.1%
Evaluating local e-e and e-n potentials 1.8%
Evaluating blip orbitals 7.9%
Updating Slater cofactor matrix 1.4%
Summing energies over MPI processes 16.0%
Charge density accumulation 1.5%
3.3 Profiling performance on GPUs
Having identified the CPU performance bottlenecks, different aspects of the simulation were prioritized in order of their respective percentages of CPU time as reported in Table 1. Vital to the success of the development cycle at this stage is repeated use of GPU profiling tools such as NVIDIA’s Nsight Systems [35] and Nsight Compute [36], and AMD’s ROCProfiler [37] and Omniperf [38]. Each of these tools is incompatible with hardware other than that of its particular vendor; NVIDIA’s Nsight ecosystem was used extensively in this work since our primary focus is on using OpenACC to port to this vendor’s GPUs.
Nsight Systems provides system-wide performance traces and metrics; a Systems profile spans the CPU, GPU, and memory, thereby facilitating the correlation of performance inhibitors across host and device. The main interface is a graphical timeline representing the database of samples collected during a given run time, and derived metrics. We typically used this timeline in the early stages of the GPU development cycle to obtain a bird’s-eye view of host-device memory copies and CUDA API calls resulting from the initial decoration of targeted code with OpenACC directives. Problematic areas were subsequently analyzed in more detail using Nsight Compute.
Nsight Compute presents detailed performance metrics obtained via hardware counters and software instrumentation, exposing the low-level GPU processes with which CUDA interfaces on a per-kernel level. It then uses a predefined set of rules to provide a guided analysis of these metrics to the user. Many of the metrics are correlated to lines of source code, and additionally to the assembly generated by the compiler; we could thus navigate directly to relevant lines in CASINO and gain insight into the mechanics underlying observed performance. These insights not only informed optimization of the directly affected source code, but also improved the efficiency of our development by preemptively mitigating GPU performance inhibitors when subsequently coding during the iterative porting process.
One of the main utilities of Nsight Compute in our development cycle was to examine GPU throughput (i.e., NVIDIA’s Speed Of Light metrics), as the achieved percentage of utilization with respect to the theoretical maximum, for both compute [i.e., the GPU’s streaming multiprocessors (SM); see Section 4 for more details] and memory; the limitations to device occupancy due to respective usage of registers and shared memory emerged as metrics of particular interest (see Section 5.1). Nsight Compute’s Launch Statistics provided details of SM and memory utilization per OpenACC parallel region, reporting the GPU resources assigned at run time to a given kernel. We used these statistics to inform experimentation with the OpenACC num_gangs, num_workers, and vector_length clauses to optimize parallel regions.
Subsequent to each optimization, a new profile was generated in order that any impact on performance could be assessed. Nsight Compute facilitates this iterative process as it can use a previously generated report as a baseline, automatically calculating any difference in metrics between this and those in the new profile.
4 Understanding and quantifying GPU performance
Modern GPUs typically comprise thousands of small, highly parallel cores and are capable of processing thousands or even millions of instructions simultaneously. The cores of a GPU are typically low-powered relative to those of a CPU, and are also designed to perform quite different tasks. Direct comparison of the performance of the former with the latter is therefore not always enlightening. Thus, to quantify better the theoretical maximum performance of the GPU it may help to take a small diversion into how GPUs work at a hardware level.
The exact naming conventions and specifics vary across different hardware vendors and even between devices from the same manufacturer. However, in general, GPUs consist of a large number of processing elements (compute clusters), each of which contains a number of GPU cores and some scheduling hardware to distribute a queue of instructions over a number of fixed-size pools of threads. Each GPU core runs the same instruction across a pool of threads in lockstep using so-called vector processors. The GPU cores themselves generally consist of a number of 16-, 32-, and 64-bit integer and floating-point vector processors and some arithmetic logic units (ALUs). Each cluster also contains a small pool of shared resources such as caches, 64- and 32-bit local registers, and some shared memory. It is ultimately these shared resources and the number of thread blocks that limit the number of compute threads the GPU can process in parallel at any one point in time.
Slightly confusingly, each device manufacturer has its own naming convention for these components, and these naming conventions are not always directly comparable. Figure 1 shows a schematic of an NVIDIA A100 GPU die and an individual SM, to help clarify the terminology.
AMD refers to its compute clusters as “compute units” and the underlying GPU cores are referred to as “stream processors.” Meanwhile NVIDIA’s terminology is more complex. They refer to compute clusters as “streaming multiprocessors.” However, they also use various marketing terms, all of which end in “core” and, crucially, none of which refer to the actual GPU cores themselves. The most commonly misused of these is “CUDA cores,” which refer to the floating-point vector processors in each cluster, not the GPU cores. They also use the term “RT cores” to refer to separate hardware used for ray-tracing and “Tensor cores” to refer to hardware for matrix operations used extensively in artificial intelligence (AI) workloads. These cores are not utilized by CASINO due to their lack of support for DP arithmetic.
(a) Diagram of a complete NVIDIA A100 GPU die [39]. The red rectangle, added by us for clarity, shows the location of a single SM.
(b) Exploded view of a single SM from an NVIDIA A100 GPU [39].
Figure 1: Schematic of an NVIDIA A100 die and an individual SM.
Another terminology worth mentioning at this stage is that used by the various programming frameworks themselves. In terms of software, GPU APIs use various terms to express the hierarchical parallelism of the underlying hardware. The exact mapping, however, is not always one-to-one, as the compiler transparently handles the mapping of the logical parallelism to the actual hardware. It is however helpful to discuss these terms as they are used in the context of different GPU programming languages.
NVIDIA’s programming language CUDA uses the terms “warps” and “thread blocks.” In simple terms a warp is the smallest compute unit the programmer has access to, a set of 32 threads, all of which work on the same program instruction in lockstep. From a hardware perspective warps can be thought of as instructions running on individual vector processors within a GPU core.
A single compute cluster can run a single instruction over multiple warps, grouped into thread blocks using the scheduling hardware. At the time of writing the current and previous two generations of NVIDIA GPUs allow each SM to handle a maximum of 64 total warps spread across a maximum of 32 thread blocks.
The CUDA API uses CUDA grid dimensions $x$ and $y$ for indexing threads. The compiler handles the mapping between these logical coordinates and the underlying hardware, so these dimensions do not necessarily need to be a multiple of 32. The compiler simply pads the instruction with zeros to map to the next largest warp. Thus the exact number of warps used for each instruction is not directly user-controllable.
AMD’s programming language ROCm helpfully uses similar terminology to CUDA for the most part, with the only differences being that “blocks” are called “workgroups” and that warps contain 64 threads.
Finally OpenACC, used by CASINO, uses similar concepts with “gangs” “workers” and “vectors”. However these in general are not directly comparable to the CUDA/ROCm terminology. “Workers” and “vectors” perform a similar role to the CUDA/ROCm $x$ and $y$ dimensions. However, the exact hardware mapping is handled transparently by the compiler. “Gangs” also do not have a direct hardware equivalent, but can be thought of as a group of threads that are scheduled together. Usually each gang is mapped to a single compute cluster but in general this is merely a suggestion from the programmer to the compiler. Note there is also no specific OpenACC term for warps; they are just blocks of 32 or 64 vectors, and programmers are generally advised to choose an appropriate number of vectors (i.e., a multiple of 32) to tune performance to whatever hardware is being used.
For these reasons, although OpenACC is hardware-agnostic by design, there are important hardware considerations to achieving good performance. In this work we are focusing particularly on using NVIDIA hardware with the NVIDIA toolkit, specifically Nsight Compute, to analyze performance. Therefore in the remainder of this work we will use the NVIDIA terminology.
To help clarify terminology we provide Table 3 to map common terms.
Table 2: Comparison of terminology used by different GPU programming languages/frameworks.
NVIDIA AMD Indicative
(CUDA) (ROCm) OpenACC mapping
SM Compute unit Gang
CudaDim.y RocmDim.y Worker
CudaDim.x RocmDim.x Vector
Warp Warp 32/64 vectors
Table 3: Terminologies used by the different vendor APIs and a typical mapping to OpenACC terms. Note that OpenACC does not guarantee any particular mapping, and this is indicative only. In particular, the compiler may assign multiple gangs to a single compute cluster. Also, an OpenACC vector is a user-defined number of threads, and often corresponds to a small number of warps, rather than just one.
5 Use of OpenACC for offloading in CASINO
5.1 Ewald interactions
To accelerate the evaluation of Ewald interactions we focused on offloading Eq. (3) to a GPU. For testing we used VMC simulations of a 3D homogeneous electron gas (HEG), with the number of particles ranging from 100 to 1630, which spans the range of particles one would use in practical simulations. This system was chosen primarily because we can easily scale the system size without having to re-optimize the wave function. The calculations were performed on an NVIDIA V100 GPU. Our initial attempt consisted of parallelizing the loop over each particle across the GPU threads in a straightforward manner. Using the OpenACC LOOP directive we computed $v_{\rm E}$ for each particle on its own thread in parallel. This in turn led to us spawning between 100 and 1630 threads on the GPU, depending on the number of particles simulated, resulting in a noticeable, if underwhelming, overall speedup of around 10–12%.
Modern GPUs can handle many hundreds of thousands of threads simultaneously without issue. Therefore, while the V100 GPU is not the latest offering from NVIDIA, we would naïvely expect significantly better performance than this from the underlying hardware. The NVIDIA V100 has 84 SMs, with a maximum of 64 warps per SM. Therefore, assuming the SMs have adequate resources, it has a theoretical maximum of 172,032 concurrent threads (that is 84 SMs running two blocks of 32 warps of 32 threads) [40].
Even with the maximum number of particles, it is clear we are not even close to saturating the GPU, so our comparatively poor performance must by caused by some other limiting factor. Commonly this is the available shared memory or the number of available registers. In our case the root cause was traced, using NVIDIA’s Nsight Compute profiler, to the GPU occupancy being limited to eight thread blocks on a single SM (512 total threads). This limit was being enforced by the GPU hardware due to the number of 64-bit GPU registers being used, which is directly proportional to the number of private variables used in the underlying loop. Thus, to improve the performance, we needed to do two things: First, reduce the number of private variables used in the loop, and hence decrease the number of GPU registers being used. Second, increase the number of available threads to make better use of the underlying hardware. This was achieved through the realization that the summations over real and reciprocal lattices used in Eq. (3) are independent of one another. Thus we can split the single, large loop over all particles into three smaller loops: one to compute the sum over real lattice vectors, one to compute the sum over reciprocal lattice vectors, and the final loop to combine them at the end to get $v_{\rm E}$. This not only reduces the number of private variables required, as the loops are smaller, but also increases the number of available threads as the loops over real and reciprocal space can be computed asynchronously. With this modification, offloading of Ewald interactions leads to an overall speedup compared to the CPU in simulations with more than 350 particles. Simulations with fewer particles give a less significant overall speedup, which is likely due to the lack of work for the GPU; fewer particles means there is less scope to use the large number of available threads. We also note that the fraction of run-time spent in the Ewald function decreases linearly with particle count, at least for large enough simulations (typically 300 particles or more). However, since the Ewald interactions only account for 36% of the overall runtime in the CPU-only case, the overall effect of GPU acceleration quickly provides diminishing returns, with the overall speedup appearing to plateau at around 1.43 times. The full results are shown in Fig. 2.
(a) Total runtime vs. number of particles.
(b) Speedup relative to the CPU vs. number of particles.
Figure 2: Runtime comparison for VMC simulations of the 3D HEG with and without offloading of Ewald interactions to the GPU. All runs were performed using all 32 cores of a POWER9 CPU @ 2.7 GHz with a single NVIDIA V100 GPU.
5.2 Long-range two-body Jastrow terms
For the long-range Jastrow factor discussed in Section 2.3 we focused our efforts on offloading the evaluation of the values and derivatives of the two-body terms $u(r)$ and $p({\bf r})$ as given by Eqs. (6) and (7), respectively. As in Section 5.1 we tested our changes in simulations of 3D HEGs with 100 to 1930 particles performed on an NVIDIA V100 GPU. This spans the range of particles one would use for practical simulations and was chosen primarily because we can easily scale the system size without having to redo the wave-function optimization. The Jastrow $u$ and $p$ terms contain random parameters, with the cutoff length for $u$ set to the radius of the largest sphere that can be inscribed in the Wigner-Seitz cell.
For both terms our initial attempts led to a significant slowdown of around 3.84 times for the $u$ term and 1.28 times for the $p$ term with 678 particles. A full breakdown of speedup for each of the terms can be found in Table 4. We investigated the cause of the slowdown for both terms using the NVIDIA profiling tools Nsight Systems and Nsight Toolkit.
Table 4: Runtimes and speedups for offloading the $p$ and $u$ two-body Jastrow terms in 3D HEG VMC calculations with 678 particles.
Activity Run type Runtime (s) Speedup %age runtime
vs. CPU
Total runtime CPU 9311.46 1.00 100%
Total runtime GPU (init. attempt) 12816.79 0.73 100%
Total runtime GPU (final version) 4307.02 2.16 100%
$P$ evaluation CPU 8513.60 1.00 91.43%
$P$ evaluation GPU (init. attempt) 10895.20 0.78 85.01%
$P$ evaluation GPU (final version) 3520.12 2.42 81.73%
$U$ evaluation CPU 114.95 1.00 1.23%
$U$ evaluation GPU (init. attempt) 442.07 0.26 5.65%
$U$ evaluation GPU (final version) 111.12 1.03 2.58%
For the $u$ term the occupancy was limited to only four warps on a single V100 SM (for OpenACC vector length of 128, i.e., 128 concurrent threads). This was similar to the situation for Ewald interactions described in Section 5.1 in that it is a hardware-imposed limit due to the number of 64-bit (DP) GPU registers required (which is directly proportional to the number of private variables used in the underlying loop). However, unlike the Ewald interactions, our options for improving this situation are much more limited. We cannot break the underlying loop down any further, as each loop iteration is its own distinct task. The only option available was to reduce the number of 64-bit variables used in the loop, for example by combining $x$, $y$, and $z$ components into three-element arrays. This somewhat improved the situation. However, the GPU performance is now merely on par with the CPU. Thus, to realize any more performance, the only option remaining is to move to using SP arithmetic, which in effect doubles the number of available registers. Therefore, we have decided to keep the $u$ term on the CPU for the production version of the code when using DP arithmetic. Meanwhile we defer the discussion of SP performance until Section 6.
For the $p$ term, the cause of the slowdown was somewhat more complex. Our initial investigations revealed that the GPU hardware was limiting the performance to a single warp on a single SM (32 threads). Further investigations then identified this as an issue with the amount of shared memory used by each GPU thread. To understand the problem we must first briefly discuss the memory hierarchy for GPUs. GPU threads generally have access to four levels of memory. NVIDIA names these as L0, L1, L2, and global memory. These can be thought of as being analogous to L1 cache, L2 cache, and DRAM for a CPU. They can be seen as light blue rectangles in Fig. 1. L0 is tiny and is used only to store the current instruction being worked on by the thread block. The L1 cache is shared across all thread blocks within a single SM and is the fastest for threads to access. However it is not very large, totalling 128 kB per SM for an NVIDIA V100 [40], and is generally only intended for frequently accessed local and temporary variables. The L2 cache is slower to access but is shared between all the SMs and is slightly larger at 6 MB [40]. Finally the global memory is the slowest to access, but can be accessed by any thread across the entire GPU and is orders of magnitude larger at 32 GB [40]. When evaluating the $p$ term, each thread needs access to three complex arrays $e^{i{\bf b}_{1}\cdot{\bf r}_{ij}}$, $e^{i{\bf b}_{2}\cdot{\bf r}_{ij}}$, and $e^{i{\bf b}_{3}\cdot{\bf r}_{ij}}$ used in Eq. (8) to compute all the required cosines for Eq. (7). These arrays were marked as private to each thread and as such had to be stored in the L1 cache. Individually the arrays are reasonably small, however, due to the large number of threads, the arrays do not all fit into the L1 or L2 caches. This was further compounded by OpenACC not allowing for direct mapping of threads to the hardware, as this is handled by the compiler. In our case the compiler was only allocating a single SM for all the threads in the loop, since the V100 can theoretically handle up to 2048 threads per SM (amounting to 32 blocks of 64 threads) [40]. This was ultimately the cause of the slowdown. Each thread block was using too much shared memory. As a result the SM was only able to schedule one thread block at a time. This was severely hampering the performance of the GPU as it was not only limiting the number of parallel threads but also wasting many CPU/GPU cycles needlessly copying data to and from shared memory. In order to fix this, we need to keep these arrays in the global device memory to allow the hardware to launch many more concurrent threads. This was achieved by allocating three large $3\times N$ 2D arrays. Each thread is then given one row of each array to work on, corresponding to the index of the current particle. This modification leads to a substantial performance increase of 2.42 times compared to the CPU for 678 particles (see Table 4) and hence a decrease in the overall runtime of around 53.74%.
The full results for offloading both the Ewald interactions and Jastrow $p$ term for 100 to 1920 particles can be found in Fig. 3.
(a) Runtime vs. number of particles.
(b) Speedup of different activities relative to the CPU vs. number of particles.
Figure 3: Runtime comparison for VMC simulations of a 3D HEG with and without offloading of the two-body Jastrow $p$ terms to the GPU. The results have been averaged over five runs. All runs were performed using all cores of a 32-core POWER9 CPU @ 2.7 GHz with an NVIDIA V100 GPU.
From this we can see that the CPU and GPU runtimes steadily diverge from one another such that by around 1200 particles the GPU version is roughly 2.4 times faster. The total speedup also appears to level off as we further increase the particle count. This is likely a hardware-imposed bottleneck as the memory usage increases linearly with particle count. Therefore as particle count increases the number of threads that can be run concurrently decreases.
One final thing to note at this stage is that the relative speedup data (Fig. 3b) appear quite noisy as a function of system size. This is due to the intrinsic variance in the runtime arising from different random walks being taken (e.g., different numbers of moves being accepted or rejected) in a short simulation; averaging over multiple runs with different random seeds at each system size allows one to reduce the apparent noise.
5.3 Evaluation of blip orbitals
Since the calculations used for the evaluation of blip orbitals consist of straightforward linear algebra, GPU offloading is likewise straightforward. The approach on the CPU makes use of the Level 1 and Level 2 basic linear algebra subprograms (BLAS) libraries, specifically: ddot, for performing the dot product of two DP vectors, and dgemv, for calculating the DP matrix-vector operation ${\bf y}\mapsto\alpha A{\bf x}+\beta{\bf y}$, where $\alpha$ and $\beta$ are scalars, ${\bf x}$ and ${\bf y}$ are vectors, and $A$ is a matrix [41].
There exist equivalent libraries for offloading BLAS calculations to a GPU. The most popular of these is cuBLAS, which is supplied by NVIDIA as part of its CUDA software suite, and is designed to be a simple drop-in replacement [13].
To test the GPU performance, we simply swapped all the calls to BLAS with their cuBLAS equivalents, while handling the data movement from CPU to GPU using OpenACC. We then performed simulations of 3D HEGs ranging from 100 to 1930 particles on an NVIDIA V100 GPU. These tests, however, proved fruitless, causing a slowdown of around five times.
Such results however, are not entirely surprising. The CPU implementation of the BLAS libraries are already highly optimized. Thus, to overcome the overhead of the data transfer from CPU to GPU, we ideally would need to use vectors with a very large number of elements. Our own benchmarks suggest the need for around 200,000 elements before breaking even when using cuBLAS for ddot (see Fig. 4 for details). CASINO blip-orbital evaluations make use of many small calls to BLAS, generally using 64 element vectors and matrices. These are not well suited for cuBLAS due to the number of data transfers from CPU to GPU and the relatively low number of threads utilized; hence the relatively poor GPU performance.
There is however still some scope for batching the calls to cuBLAS to be run concurrently using CUDA streams, thus allowing us to better utilize the GPU hardware by calculating several functions in parallel. This approach however would require significantly more work and as such has been left for future investigation.
(a) Runtime vs. number of vector elements.
(b) Speedup relative to the CPU vs. number of vector elements.
Figure 4: Runtime comparison of OpenBLAS (on CPU) vs. cuBLAS (on GPU) for the ddot function with an increasing number of vector elements. These calculations were performed using all cores of a 32-core POWER9 CPU @ 2.7 GHz and an NVIDIA V100 GPU.
6 Use of SP arithmetic in CASINO
6.1 Switching precision
The majority of real and complex variables in CASINO are of DP kind. The only exceptions are that (i) SP variables are used in some low-level tasks such as timing routines, and (ii) blip coefficients can optionally be stored as SP numbers, to reduce memory requirements. Adopting SP throughout CASINO offers a significant speedup on a CPU and potentially an even more significant speedup on a GPU.
Due to the use of Fortran, it is straightforward to switch precision. The “kind” of real and complex numbers used throughout CASINO is a parameter defined in a single module. A small BASH script (i) changes the default kind parameter to SP and (ii) replaces calls to DP BLAS and LAPACK routines with calls to their SP equivalents. The kind parameter is used to select the appropriate kind flag for calls to MPI routines. A few parts of the code relating to the initial setup of a calculation, such as the evaluation of lookup tables for Ewald interactions, are numerically sensitive and are always performed using DP arithmetic. Substituting the BLAS and LAPACK routines is easy because of the rigorous naming conventions for these routines. The apparently obvious alternative of overloading external BLAS and LAPACK routines is nontrivial because of the use of assumed-size arrays in these routines. For other external libraries used by CASINO, such as Netlib’s specfun library [42] for special functions, we simply overload wrapper functions in CASINO.
To enable the successful use of SP arithmetic it was necessary to re-express many hardwired tolerance parameters throughout the CASINO software in terms of the machine precision (and in one or two places in terms of the largest and smallest representable numbers). Fortunately Fortran provides intrinsic functions for just this purpose.
6.2 Accuracy of SP arithmetic in QMC calculations
Switching from DP to SP makes surprisingly little difference to VMC and DMC results when a blip basis set is used for the orbitals. For example, the VMC total energy of a $2\times 2\times 2$ supercell of bulk Si at the simulation-cell Baldereschi point using Trail-Needs Si pseudopotentials was found to be $-7.875606(14)$ Ha per primitive cell with SP arithmetic and $-7.875615(10)$ Ha per primitive cell with DP arithmetic. In these calculations, a Slater-Jastrow trial wave function was used, with $u$, $\chi$, $f$, and $p$ terms in the Jastrow exponent and complex Perdew-Burke-Ernzerhof orbitals represented in a blip basis set in the Slater determinant. The parameters in the Jastrow factor were optimized by energy minimization with DP arithmetic. The corresponding DMC energies at a time step of 0.04 Ha-1 were found to be $-7.88720(4)$ and $-7.88713(4)$ Ha per primitive cell using SP and DP arithmetic, respectively. Similarly excellent agreement between SP and DP QMC results is found in a $2\times 2\times 2$ supercell of Si at the simulation-cell $\Gamma$ point, where the trial wave function is real. These results demonstrate that SP arithmetic is perfectly adequate for at least some QMC calculations.
On the other hand, when an optimized Gaussian basis set is used for a $2\times 2\times 2$ supercell of Si at the $\Gamma$ point with a Troullier-Martins Si pseudopotential, the Slater-Jastrow VMC SP and DP energies are $-7.88573(3)$ and $-7.88283(3)$ Ha per primitive cell, respectively. These differ significantly, indicating that Gaussian basis sets are more challenging for SP arithmetic. It was verified that this difference is not affected by reasonable changes to the tolerance controlling the distances beyond which Gaussian basis functions are truncated to zero. The difference in SP and DP total energies is equal to the corresponding difference in kinetic energies, while the potential energies are in statistical agreement, implying that the evaluation of the Laplacians of orbitals expanded in a Gaussian basis set is the part that is particularly sensitive to the choice of precision.
Furthermore, even when a blip basis set is used, wave-function optimization is less effective with SP arithmetic than with DP arithmetic. Wave function optimization is usually computationally cheaper than DMC, but is the most technically challenging aspect of QMC work and the most costly in terms of human time. Re-optimization by energy minimization with SP arithmetic does not significantly alter or spoil the previously optimized trial wave function used in the results quoted above for a $2\times 2\times 2$ supercell of Si at the Baldereschi point. However, optimization of the parameters in that trial wave function from scratch with SP arithmetic yields a VMC energy of $-7.86165(9)$ Ha per primitive cell, which is significantly higher than the VMC energy of $-7.875606(14)$ Ha per primitive cell obtained using energy minimization with DP arithmetic.
In summary, although SP arithmetic can safely be used for many types of QMC calculation, there exist important use-cases in which SP arithmetic is problematic. In the present work we are primarily concerned with the speedup that can be obtained through the use of SP arithmetic; we will not investigate the accuracy of such calculations further.
6.3 Speedups due to offloading with SP arithmetic
Table 5 shows the speedup resulting from offloading a 3D HEG VMC simulation using SP arithmetic for 678 and 1290 particles respectively. These have been broken down into contributions from the Ewald interactions as well as the Jastrow $u$ and $p$ terms. From this we can see that both the Ewald interactions and Jastrow $p$ term show a similar performance improvement to the DP case.
Table 5: Runtimes and speedups for offloading the $p$ and $u$ two-body Jastrow terms and Ewald interactions in VMC simulations of 3D HEGs with $N=678$ and 1290 particles. SP arithmetic was used.
Activity Run type $N$ Runtime (s) Speedup vs. CPU
Total runtime CPU 678 876.90 1.00
Total runtime GPU 678 346.71 2.53
$P$ eval. CPU 678 545.44 1.00
$P$ eval. GPU 678 193.11 2.82
$U$ eval. CPU 678 4.31 1.00
$U$ eval. GPU 678 3.88 1.11
Ewald eval. CPU 678 20.79 1.00
Ewald eval. GPU 678 11.63 1.79
Total runtime CPU 1290 3225.20 1.00
Total runtime GPU 1290 958.58 3.36
$P$ eval. CPU 1290 1958.12 1.00
$P$ eval. GPU 1290 620.70 3.15
$U$ eval. CPU 1290 8.75 1.00
$U$ eval. GPU 1290 7.65 1.14
Ewald CPU 1290 78.10 1.00
Ewald GPU 1290 21.53 3.63
From this we can see that both the Ewald interactions and Jastrow $p$ term see a similar performance improvement as the DP case, leading to speedups of 2.82 and 1.79 times, respectively, for 678 and 1290 particles.
Unfortunately, despite doubling the number of available registers, the $u$ term still sees only a small improvement of 1.11 and 1.14 times respectively. Further investigations revealed that the GPU kernel used in the computation was experiencing significant “warp stall,” meaning that the SM is waiting multiple cycles for the current instruction to complete. This is indicative that there is simply too much work being performed per thread. Ideally for best performance on a GPU we want many threads each performing a small amount of work. This however, is not the case for the $u$ term as each iteration of the corresponding loop performs significant work which cannot easily be broken down into smaller parallel tasks. As such, we conclude that offloading the $u$ term to GPU is simply not worthwhile with present hardware and this has not been enabled in the release version of the code. However, given the industry trend towards AI (which heavily favors the use of SP and even half-precision arithmetic) a significant performance improvement for SP workloads is not out of the question, so this may be worth revisiting in future with newer hardware.
The full results for offloading both the Ewald interactions and Jastrow $p$ term for 100 to 1920 particles can be found in Fig. 5. For the CPU case we can see that the use of SP arithmetic leads to a performance improvement of around 1.5 times. This is broadly in line with what we would expect from the POWER9 hardware used, with the smaller datatype allowing for improved CPU cache performance, reduced memory bandwidth, and more efficient CPU vectorization [43]. The SP GPU case also sees a significant performance improvement, being between 1.9 and 6 times faster than the DP CPU version, and between 1.5 and 1.8 times faster than the DP GPU case. GPUs are primarily intended for work with computer graphics, where operations involving SP and even half-precision are commonplace; indeed on consumer cards and some cheaper server cards it is common to have more SP floating-point units than DP ones by a ratio of as much as 30:1. Therefore the large performance increases we see here are unsurprising as GPU hardware has been designed and optimized for performance with SP arithmetic.
(a) Runtime vs. number of particles.
(b) Speedup relative to DP CPU vs. number of particles.
(c) Speedup relative to DP GPU vs. number of particles.
Figure 5: Runtime comparison for VMC simulations of a 3D HEG with DP and SP arithmetic averaged over five runs. All runs were performed using all cores of a 32-core POWER9 CPU @ 2.7 GHz with an NVIDIA V100 GPU.
7 Performance on an NVIDIA H100
So far we have only discussed performance on an NVIDIA V100 GPU. This particular GPU, whilst still performant, is now somewhat dated, having been released in 2017. Therefore we have also tested the performance of CASINO on the NVIDIA H100 GPU. This GPU is based on the NVIDIA Hopper architecture and is two generations ahead of the V100. The H100 is expected to be significantly faster than the V100, with a reported 2.5 times increase in peak performance for DP operations [40, 44].
To investigate the performance of the H100, we used the NVIDIA Grace-Hopper GH200 Superchip allocation on Bede. This consists of a 72 core NVIDIA Grace 64-bit ARM-based CPU @ 3.483 GHz with a 96 GB NVIDIA H100 GPU.
For our tests we have studied the 3D HEG with 118 to 1640 particles and DP arithmetic. The results of these tests can be seen in Fig. 6.
We note at this stage that a direct comparison of performance vs. the CPU is essentially meaningless as the CPU used in the Grace-Hopper system is an ARM-based CPU, which is not designed for compute-intensive workloads and as such is not performant for QMC calculations.
Hence we will only compare the performance of the H100 with the V100 for the sections of code which we have accelerated using the GPU, i.e., Ewald interactions and the Jastrow $p$ term.
From this we can see that for both the Ewald interactions and Jastrow $p$ term the H100 is significantly faster. For the $p$ term the H100 is around 1.6 times faster for 118 particles and the speedup relative to the V100 slowly rises with particle count to around 2.1 times faster for 1642 particles. However, for the Ewald interactions the H100 is 2.4 times faster than the V100 for 118 particles, after which the speedup rises and then peaks at 2.55 times faster for 226 particles, before slowly decreasing to 1.61 times at 1640 particles. All these performance increases are in line with the expected performance increase for DP operations.
The initially higher performance for Ewald interactions at low particle counts is likely due to the lack of available work for the GPU. As the number of particles increases, so does the number of threads, and thus the kernel launch time. For the V100 this effect was mostly obscured by the data transfer time, which is in general an order of magnitude larger than the kernel launch time.
However, in the case of the H100, the Grace-Hopper superchip grants both the CPU and the GPU direct access to each other’s memory space. Therefore the data transfer times are essentially instantaneous, and the effect of the increased kernel launch times is much more readily visible, with the 2.4 to 2.55 times speedup being essentially in line with Nvidia’s own estimates for the theoretical maximum performance uplift of 2.6 times [39].
(a) Runtime vs. number of particles.
(b) Speedup of GPU accelerated sections relative NVIDIA V100 vs. number of particles.
Figure 6: Runtime comparison for VMC simulations of a 3D HEG. Runs were performed using a 32-core POWER9 CPU @ 2.7 GHz paired with an NVIDIA V100 GPU and an NVIDIA Grace 72-core arm64 CPU @ 3.483 GHz paired with an NVIDIA H100 GPU.
8 Further examples and discussion
8.1 2D HEG code performance
CASINO is a complex codebase and, while the 3D HEG is useful for testing and gauging the general offloading performance, it is lacking in two main areas. Firstly it is a 3D system, so it does not show how GPU offloading affects 2D systems, which are a significant use-case and which use a slightly different code path due to various simplifications and optimizations. Secondly and more importantly, the HEG does not include the complications that arise from the presence of atomic cores in ab initio QMC calculations, such as nontrivial single-particle orbitals, pseudopotentials, and inhomogeneous Jastrow terms. These more complex calculations are less dependent on homogeneous two-body Jastrow factors and Ewald interactions. We therefore study two final examples: a 2D HEG, to demonstrate 2D code performance; and a $9\times 9$ supercell of monolayer hexagonal boron nitride (hBN), to demonstrate the performance for a more realistic use-case of CASINO.
For our 2D test we use an equivalent example to the 3D case: a 2D HEG with between 100 and 1642 particles. The calculations were again performed on a POWER9 CPU with an NVIDIA V100 GPU and were averaged over five runs. The results of these calculations can be see in Fig. 7. Unlike the 3D case the 2D starts out fractionally slower on the GPU, likely due to the low number of threads. However, this quickly changes as the system size increases, with the GPU version being around 3.5 times faster for 1200 particles. The speedup for the 2D HEG appears to level off with system size, as in the 3D case, although noticeably more slowly. The speedup for the Ewald interactions increases at a much lower rate than in the 3D case, and appears less linear with system size. We also note that offloading the Ewald interaction at 98 particles appears to show a small speedup of 1.1 times. In reality, however, this is an artifact of the runtimes being small and very similar to one another. For the CPU version, the Ewald interactions took an average of 0.21 seconds with a standard deviation of 0.04 seconds, compared with 0.19 seconds with a standard deviation of 0.03 seconds for the GPU version. This means the values were within one standard deviation of each other, and as such this “phantom” speedup can be regarded as statistically insignificant.
(a) Runtime vs. number of particles.
(b) Speedup relative to the CPU vs. number of particles.
Figure 7: Runtime comparison for VMC simulations of a 2D HEG, with and without offloading of the two-body Jastrow $p$ term to the GPU, averaged over five runs. All runs were performed using all cores of a 32-core POWER9 CPU @ 2.7 GHz with an NVIDIA V100 GPU.
8.2 Hexagonal boron nitride
Our second example is a $9\times 9\times 1$ supercell of hexagonal boron nitride.
(a) Total runtime and times spent computing Ewald interactions and the Jastrow $p$ term.
(b) Speedup relative to the CPU.
Figure 8: Runtime comparison for a VMC simulation of a $9\times 9$ supercell of monolayer hBN. All runs were performed using all cores of a 32-core POWER9 CPU @ 2.7 GHz with an NVIDIA V100 GPU.
Figure 8 shows the total runtime of the simulation, alongside the time spent computing the Ewald interactions and the Jastrow $p$ term. From this we can see that, while we do see significant speedups in the computation of both the Ewald interactions and the Jastrow $p$ term, we do not see as significant an improvement for the overall runtime, because both sections take up a smaller fraction of the overall runtime than they do in the HEG.
9 Conclusions
We have investigated the speedup to the CASINO code that can result from using OpenACC to offload various aspects of a quantum Monte Carlo calculation with a Slater-Jastrow trial wave function to a GPU. For electron gases in two and three dimensions we find that, on a POWER9 CPU with an NVIDIA V100 GPU, offloading the evaluation of Ewald interactions is beneficial when more than about 350 electrons are present; with 1000 electrons, offloading accelerates the evaluation of Ewald interactions by a factor of $\sim 2.5$. Likewise, offloading the computation of long-range two-body terms in the Jastrow correlation factor is beneficial with more than about 150 electrons, speeding up Jastrow factor evaluation by a factor of (again) $\sim 2.5$ with 1000 electrons. The use of SP arithmetic increases the speedup due to offloading to a factor of about three (on top of the speedup of about 1.3 to the CPU version due to the use of SP). In electron gases, the evaluation of two-body Jastrow terms and Ewald interactions account for a significant fraction of the overall runtime, and so these accelerations translate into a similar speedup for the entire calculation.
For ab initio QMC calculations, the speedups to Ewald interactions and two-body Jastrow terms are in line with the above results for electron gases, but there are many more contributing factors to the total runtime, including inhomogeneous Jastrow correlations, the evaluation of single-particle orbitals in a localized basis set, and the evaluation of pseudopotentials. For a $9\times 9$ supercell of monolayer hexagonal BN we find an overall speedup of 1.17 due to offloading.
These results were obtained by comparing timings for a single process with a single thread running on a single CPU, with and without offloading to a single GPU. An important issue to investigate is the optimal use of GPUs on compute nodes with multiple MPI processes competing for the same set of GPUs. Furthermore it would be desirable to investigate the benefits of offloading on a range of different hardware. Finally, this work has focussed on the kinds of QMC problem of interest in condensed-matter physics: we have studied electron gas models and a 2D semiconductor. In large molecular systems, electron-electron-nucleus ($f$) Jastrow terms are relatively important, and an obvious future development would be to investigate the offloading of these Jastrow terms.
Acknowledgements
This work was funded by EPSRC (Grant No. EP/W026775/1). The work made use of the facilities of the N8 Centre of Excellence in Computationally Intensive Research (N8 CIR) provided and funded by the N8 research partnership and EPSRC (Grant No. EP/T022167/1). The Centre is coordinated by the Universities of Durham, Manchester, and York. Computer resources were also provided by Lancaster University’s High End Computing cluster. We acknowledge useful conversations with M. Pacey.

### v2

@@
 In these methods the quantum many-body problem is tackled directly in the continuous, $3N$-dimensional position basis, with Monte Carlo methods being used to explore and integrate over the high-dimensional configuration space.
-In this article we focus on the variational and diffusion quantum Monte Carlo (VMC and DMC) methods, which allow one to find the ground state and selected excited states of fermionic and bosonic systems, as implemented in the CASINO software [1].
+In this article we focus on the variational and diffusion quantum Monte Carlo (VMC and DMC) methods, which allow one to find the ground state and selected excited states of fermionic and bosonic systems, as implemented in the CASINO software [1, 2].
 A major challenge for these methods is the computational expense, which not only constrains the range of systems that can be studied, but also limits the accuracy that can be achieved, as users are forced to use smaller simulation supercells, larger time steps, etc.
@@
 The OpenACC programming standard and API [6] is supported by Fortran compilers including NVIDIA’s nvfortran, the Cray/HPE Fortran compiler, and GNU’s gfortran.
-OpenACC is developed and overseen by the OpenACC Organization, a non-profit organization whose members include the NVIDIA Corporation, AMD, HPE, ARM, and Oak Ridge National Laboratory (home to the Frontier supercomputer), as well as many HPC centers and academic institutions.
+OpenACC is developed and overseen by the OpenACC Organization, a non-profit organization whose members include the NVIDIA Corporation, AMD, HPE, ARM, and Oak Ridge National Laboratory (home to the Frontier supercomputer), as well as many high-performance computing centers and academic institutions.
+OpenACC fully supports Fortran, C, and C++.
 OpenACC compiler directives are used to identify parts of a program that should be compiled for execution on an accelerator device; furthermore these directives are also used to provide much of the functionality specific to programming for GPUs.
@@
 (3) [23].
-Similar expressions can be derived for Ewald interactions in two dimensions [23, 24, 25].
-It is clear that the cost of evaluating the Ewald interaction in Eq.
-(3) is vastly greater than the cost of evaluating $1/r$.
-Equation (3) is suitable for offloading to GPUs because the evaluation of the series is dominated by floating point operations and both the real- and reciprocal-space series are inherently parallelizable; moreover, much of the data required can be moved to the GPU once, at the outset of the calculation.
-Offloading of Eq.
-(3) is investigated in Section 5.1.
+Similar expressions can be derived for Ewald interactions in two dimensions [24, 25, 26]:
+$v_{\rm E}({\bf r})$ $=$ $\frac{\pi}{A_{\rm s}}\sum_{{\bf G}_{\rm s}\neq{\bf 0}}\frac{e^{zG_{\rm s}}\operatorname{erfc}\left(\frac{G_{\rm s}}{2\sqrt{\gamma}}+z\sqrt{\gamma}\right)+e^{-zG_{\rm s}}\operatorname{erfc}\left(\frac{G_{\rm s}}{2\sqrt{\gamma}}-z\sqrt{\gamma}\right)}{G_{\rm s}}e^{i{\bf G}_{\rm s}\cdot{\bf r}}$ (5)
+${}+\sum_{{\bf R}_{\rm s}}\frac{\operatorname{erfc}\big(\sqrt{\gamma}|{\bf r}-{\bf R}_{\rm s}|\big)}{|{\bf r}-{\bf R}_{\rm s}|}-\frac{2\pi}{A_{\rm s}}\left[z\operatorname{erf}(z\sqrt{\gamma})+\frac{e^{-\gamma z^{2}}}{{(\gamma\pi)}^{\frac{1}{2}}}\right]$
+$v_{\rm M}$ $=$ $\frac{2\pi}{A_{\rm s}}\sum_{{\bf G}_{\rm s}\neq{\bf 0}}\frac{\operatorname{erfc}\left(\frac{G_{\rm s}}{2\sqrt{\gamma}}\right)}{G_{\rm s}}+\sum_{{\bf R}_{\rm s}\neq{\bf 0}}\frac{\operatorname{erfc}\left(\sqrt{\gamma}R_{\rm s}\right)}{R_{\rm s}}-\frac{2\sqrt{\gamma}}{\pi^{\frac{1}{2}}}-\frac{2\pi^{\frac{1}{2}}}{A_{\rm s}\sqrt{\gamma}},$ (6)
+where $A_{\rm s}$ is the area of the simulation cell.
+It is clear that the cost of evaluating the Ewald interaction in Eqs.
+(3) or (5) is vastly greater than the cost of evaluating $1/r$.
+Equations (3) and (5) are suitable for offloading to GPUs because the evaluation of the series is dominated by floating point operations and both the real- and reciprocal-space series are inherently parallelizable; moreover, much of the data required can be moved to the GPU once, at the outset of the calculation.
+Offloading of Ewald interactions is investigated in Section 5.1.
 2.3 Long-range Jastrow terms
@@
 $p({\bf r})$ $=$ $\sum_{m=1}^{N_{p}}a_{m}\sum_{{\bf G}_{\rm s}\in\bigstar_{m}^{+}}\cos\left({\bf G}_{\rm s}\cdot{\bf r}\right),$ (9)
-[…]
+where $\{\alpha_{i}\}$ are optimizable parameters for $i=0$ and $i=2,\ldots,N_{u}$, while $\alpha_{1}$ is determined by the Kato cusp conditions, $L_{u}$ is a cutoff length that must be less than half the distance from the origin to the closest nonzero simulation-cell lattice point, $\Theta$ is the Heaviside function and $C$ is the number of derivatives of $u(r)$ that are well-defined at $r=L_{u}$; usually we have $C=2$ or $C=3$.
+Typically the expansion order $N_{u}$ is chosen to be in the range 8–12.
+The $\{a_{m}\}$ are optimizable coefficients for each star $\bigstar_{m}$ of simulation-cell reciprocal-lattice points ${\bf G}_{\rm s}$; the “+” superscript indicates that only one ${\bf G}_{\rm s}$ vector out of each $\pm{\bf G}_{\rm s}$ pair is included in the sum.
+Typically $N_{p}=3$1515 stars of ${\bf G}_{\rm s}$ vectors are used.
 For ${\bf G}_{\rm s}=n_{1}{\bf b}_{1}+n_{2}{\bf b}_{2}+n_{3}{\bf b}_{3}$, where $\{{\bf b}_{i}\}$ are supercell reciprocal lattice vectors, we may write
@@
 The powers of these exponentials are computed and buffered.
-The required number of reciprocal lattice points is typically $\sim$10$\sim$100.
+The required number of reciprocal lattice points is typically $\sim$$\sim10$0.
 We note that for large plane-wave expansions there is scope to reduce the computational cost of the $p$ term by means of a blip (B-spline) re-representation.
@@
 Switching from DP to SP makes surprisingly little difference to VMC and DMC results when a blip basis set is used for the orbitals.
-For example, the VMC total energy of a $2\times 2\times 2$ supercell of bulk Si at the simulation-cell Baldereschi point using Trail-Needs Si pseudopotentials was found to be $-7.875606(14)$ Ha per primitive cell with SP arithmetic and $-7.875615(10)$ Ha per primitive cell with DP arithmetic.
+For example, the VMC total energy of a $2\times 2\times 2$ supercell of bulk Si at the simulation-cell Baldereschi point using Trail-Needs Si pseudopotentials was found to be 7.875606(14)-7.875606(14) Ha per primitive cell with SP arithmetic and 7.875615(10)-7.875615(10) Ha per primitive cell with DP arithmetic.
 In these calculations, a Slater-Jastrow trial wave function was used, with $u$, $\chi$, $f$, and $p$ terms in the Jastrow exponent and complex Perdew-Burke-Ernzerhof orbitals represented in a blip basis set in the Slater determinant.
 The parameters in the Jastrow factor were optimized by energy minimization with DP arithmetic.
-The corresponding DMC energies at a time step of 0.04 Ha-1 were found to be $-7.88720(4)$ and $-7.88713(4)$ Ha per primitive cell using SP and DP arithmetic, respectively.
+The corresponding DMC energies at a time step of 0.04 Ha-1 were found to be 7.88720(4)-7.88720(4) and 7.88713(4)-7.88713(4) Ha per primitive cell using SP and DP arithmetic, respectively.
 Similarly excellent agreement between SP and DP QMC results is found in a $2\times 2\times 2$ supercell of Si at the simulation-cell $\Gamma$ point, where the trial wave function is real.
 These results demonstrate that SP arithmetic is perfectly adequate for at least some QMC calculations.
-On the other hand, when an optimized Gaussian basis set is used for a $2\times 2\times 2$ supercell of Si at the $\Gamma$ point with a Troullier-Martins Si pseudopotential, the Slater-Jastrow VMC SP and DP energies are $-7.88573(3)$ and $-7.88283(3)$ Ha per primitive cell, respectively.
+On the other hand, when an optimized Gaussian basis set is used for a $2\times 2\times 2$ supercell of Si at the $\Gamma$ point with a Troullier-Martins Si pseudopotential, the Slater-Jastrow VMC SP and DP energies are 7.88573(3)-7.88573(3) and 7.88283(3)-7.88283(3) Ha per primitive cell, respectively.
 These differ significantly, indicating that Gaussian basis sets are more challenging for SP arithmetic.
@@
 Re-optimization by energy minimization with SP arithmetic does not significantly alter or spoil the previously optimized trial wave function used in the results quoted above for a $2\times 2\times 2$ supercell of Si at the Baldereschi point.
-However, optimization of the parameters in that trial wave function from scratch with SP arithmetic yields a VMC energy of $-7.86165(9)$ Ha per primitive cell, which is significantly higher than the VMC energy of $-7.875606(14)$ Ha per primitive cell obtained using energy minimization with DP arithmetic.
+However, optimization of the parameters in that trial wave function from scratch with SP arithmetic yields a VMC energy of 7.86165(9)-7.86165(9) Ha per primitive cell, which is significantly higher than the VMC energy of 7.875606(14)-7.875606(14) Ha per primitive cell obtained using energy minimization with DP arithmetic.
 In summary, although SP arithmetic can safely be used for many types of QMC calculation, there exist important use-cases in which SP arithmetic is problematic.
@@
 Table 5: Runtimes and speedups for offloading the $p$ and $u$ two-body Jastrow terms and Ewald interactions in VMC simulations of 3D HEGs with $N=678$ and 1290 particles. SP arithmetic was used.
-Activity Run type $N$ Runtime (s) Speedup vs. CPU
+Activity Run type NN Runtime (s) Speedup vs. CPU
 Total runtime CPU 678 876.90 1.00
@@
 In electron gases, the evaluation of two-body Jastrow terms and Ewald interactions account for a significant fraction of the overall runtime, and so these accelerations translate into a similar speedup for the entire calculation.
+A speedup of $2.5$ reduces random errors on a QMC calculation by a factor of $1/\sqrt{2.5}\approx 0.63$ for a given computational resource.
+Or, for a given error bar on the total energy and a given computational resource, CASINO could be used to study a system that is $2.5^{1/3}\approx 1.4$ times larger.
 For ab initio QMC calculations, the speedups to Ewald interactions and two-body Jastrow terms are in line with the above results for electron gases, but there are many more contributing factors to the total runtime, including inhomogeneous Jastrow correlations, the evaluation of single-particle orbitals in a localized basis set, and the evaluation of pseudopotentials.
@@
 These results were obtained by comparing timings for a single process with a single thread running on a single CPU, with and without offloading to a single GPU.
+Our approach of offloading two-body Jastrow terms is applicable in VMC, DMC, or reptation quantum Monte Carlo software, while offloading of Ewald interactions is also applicable in path integral Monte Carlo calculations.
 An important issue to investigate is the optimal use of GPUs on compute nodes with multiple MPI processes competing for the same set of GPUs.